OpenStudy (babbs12):

If 2x^2+y^2=-2 then evaluate d^2y/dx^2 when x=2 and y=3. Round your answer to 2 decimal places.

OpenStudy (babbs12):

@SolomonZelman can you help me again?

OpenStudy (owen3):

2x^2+y^2=-2 2(2x) + 2y * dy/dx = 0 4x + 2y dy/dx = 0 its easier if we use y ' to stand for dy/dx 4x + 2y * y ' = 0 y' = -4x /(2y) y' = -2x / y now take derivative again

OpenStudy (babbs12):

would the derivative of y be 0?

OpenStudy (owen3):

nope, derivative of y is dy/dx, or y'

OpenStudy (babbs12):

ok

OpenStudy (babbs12):

so it would be (y'*-2x)-(-2*y)y^2 right?

OpenStudy (babbs12):

(y'*-2x)-(-2*y)/y^2 is what i meant

OpenStudy (owen3):

y' = -2x / y y ' ' = ( y(-2) - (-2x)*y ' ) / y^2

OpenStudy (babbs12):

thank you...wasn't sure where to start on it

OpenStudy (owen3):

but you have to substitute back

OpenStudy (babbs12):

right

OpenStudy (owen3):

this is what you should get http://prntscr.com/9puxo8

OpenStudy (babbs12):

so... \[y"=((3(-2))-((-2(2))(-2(2)/3))/(3^2)\]

OpenStudy (owen3):

you should get -34/27 http://prntscr.com/9puynv

OpenStudy (babbs12):

y"=((-6)-(-4*(-4/3)))/9 =(-6-16/3)/9 =(-18/3-16/3)/9 =(-34/3)/9 =-34/27

OpenStudy (babbs12):

now i say thank you